In a world with 2 villages A, B and a straight river, you have a lady in village A that need to fill water in the river and next go to village B. How will she find the point, x on the river such that the time walking in straight line from A to it Ax and from it to B Bx, will be the shortest, minimising Ax +Bx?
1
Expert's answer
2020-04-23T15:17:32-0400

As per the question,

Let the co-ordinate of the village A is (A, b) and the co-ordinate of the village B is (a, B)

Let the co-ordinate of the point of on the river is O, (x, 0)

As per the question, point A and B are lying on the straight line, so slop of AO and OB will be same,

"\dfrac{B-b}{A-a}=\dfrac{-b}{A-x}"

"-b(x-a)=B(A-x)"

the equation of line passing from the point (x, 0) and and (A, b)

"y-b=\dfrac{-b}{x-A}(x-A)"

Now substituting the values,

"y-b=-\dfrac{B-b}{A-a}(x-A)"

"x=\dfrac{A-a}{b-B}(y-b)+A"

Now, taking the differenciation of both side with respect to,

"dx=\dfrac{A-a}{b-B}dy"

for maxima and minima

"\dfrac{dx}{dy}=0"

"A=a" , so distance will be minimum, if a=0 and b=0

Hence, "Ax+Bx" will be the minimum distance.